A certain material has an index of refraction of 2.0. Find (a) the speed of light in that material if light is incident from air to this material at 30o, (b) the angle between the reflected ray and the incident ray, and (c) the angle of refraction. (d) By how much is the incident light deviated?


1
Expert's answer
2021-03-31T09:32:26-0400

a)


"v=\frac{3\cdot10^8}{2}=1.5\cdot10^8\frac{m}{s}"

b)


"\alpha=2(30)=60\degree"

c)


"\beta =\arcsin{\frac{\sin{30}}{2}}=14.5\degree"

d)


"\gamma=30-14.5=15.5\degree"